OpenStudy (anonymous):

Right-endpoint sum on interval [0, 4]

OpenStudy (fibonaccichick666):

ok, upper darbaux sum way or just the calc II way?

OpenStudy (anonymous):

Can i pull this out of the sum? \[\frac{ 24n }{ n+n ^{2} }\]

OpenStudy (anonymous):

the Calc I way, whatever that is

OpenStudy (anonymous):

on the interval [0,4] I use: f(1) + f(2) + f(3) + f(4) correct?

OpenStudy (fibonaccichick666):

the bottom is imaginary?

OpenStudy (fibonaccichick666):

oh nvm, no, you can't you can pull the top out though

OpenStudy (anonymous):

i is referring to the i value for 0 < i < n

OpenStudy (fibonaccichick666):

yea, I'm used to n for that

OpenStudy (fibonaccichick666):

so anyways, you will have to factor the n on the bottom

OpenStudy (anonymous):

\[\Delta x = \frac{ b-a }{ n } = \frac{ 4-0 }{ n } = \frac{ 4 }{ n }\] I think

OpenStudy (fibonaccichick666):

you know n

OpenStudy (fibonaccichick666):

n goes to infty

OpenStudy (fibonaccichick666):

but anyways, can you pull out all the n's you possibly can and tell me what you are left with?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} 24n \sum_{i=1}^{n} \frac{ 1 }{ 4n+n^{2} } ?\]

OpenStudy (fibonaccichick666):

can you pull out any on the bottom?

OpenStudy (anonymous):

i mean 4in + 4n^2

OpenStudy (fibonaccichick666):

it's ok

OpenStudy (anonymous):

is that legal?

OpenStudy (fibonaccichick666):

what legal?

OpenStudy (fibonaccichick666):

just factor the bottom there

OpenStudy (perl):